A system has the Carnot efficiency of 16% and its cold reservoir is at 18°C. Determine the temperature of the hot reservoir in [°C].
Answer #1

As we know that for a Carnot Cycle Heat Engine :

E = [TH - TL] / [ TH ]

Here, E = 0.16, TH = ?, TL = 273+18 = 291 K

Putting all the values in the above expression -

0.16 = [TH - 291]/TH

=> 0.16*TH = TH - 291

=> 0.84*TH = 291

=> TH = 291/0.84 = 346.43 K

So temperature of hot reservoir in degree centigrade = 346.43 - 273 = 73.430C